# Evaluating a Digital Speech Therapy App for Stuttering: A Pilot Validation Study

### v1

@@ Abstract
Stuttering is a speech disorder that disrupts fluency and often leads to significant psychological and social challenges. This study evaluates the effectiveness of Eloquent, a digital speech therapy app, by analyzing pre-therapy and post-therapy speech samples using the Stuttering Severity Index-4 (SSI-4) and the S24 communication and attitude scale. Results indicate significant improvements in fluency, with reductions in SSI-4 scores across reading, speaking, duration, and physical concomitant metrics. Additionally, participants demonstrated a more positive attitude towards communication post-therapy, as evidenced by lower S24 scores. These findings highlight the potential of technology-driven, structured speech therapy interventions to deliver measurable improvements in stuttering severity and communication confidence.

### v2

@@ Abstract
Stuttering is a clinical speech disorder that disrupts fluency and leads to significant psychological and social challenges. This study evaluates the effectiveness of Eloquent, a digital speech therapy app for stuttering, by analyzing pre-therapy and post-therapy speech samples using the Stuttering Severity Index-4 (SSI-4) and the S24 communication and attitude scale. Results showed a 52.7% reduction in overall SSI-4 scores, with marked improvements in reading (45%), speaking (46%), duration (57%), and physical concomitants (63%) scores. Over 75% of participants improved by at least one severity category. S24 scores decreased by 33.5%, indicating more positive self-perceptions of speech and reduced avoidance. These findings highlight the potential of structured, technology-driven speech therapy interventions to deliver measurable improvements in stuttering severity and communication confidence.
@@ 9 Conclusion
Declaration of Competing Interests
Urvisha Shethia and Vedali Inamdar are expert consultants working with Iyaso, which is involved in the development of the digital speech therapy intervention evaluated in this study.